# Memory Mosaics at scale

### v1

@@ 5.3 In-context learning
In this section, we adopt a few-shot learning setup where each “shot” consists of one $(x,y)$ example from each possible target label category. By collecting multiple shots, we create an $n$-shot classification task. To encode these $(x,y)$ examples for memory mosaics v2 and transformers, we serialize the $(x,y)$ pairs into a sequence followed by a test query $x_{test}$. [^12] Transformers are known to be sensitive to the prompt strategies Gupta et al. (2024); Mirzadeh et al. (2024), such as the delimiter before $x$ and $y$, shuffling/not-shuffling the $(x,y)$ examples within each shot. To reduce the influence of prompt strategies, we evaluate each classification task with different delimiters (“[space]” and “\$\backslash n$”), shuffled/non-shuffled $(x,y)$ examples. Then choose the best prompt strategy for each $n$-shot classification task. Check appendix G for prompt examples. A prompt example is:
@@
Figure 3 compares the performance of memory mosaics v2 and transformers in three classification tasks with semantic target labels. The horizontal axis represents the number of shots, while the vertical axis represents the classification accuracy on $x_{test}$. We can observe two phenomena: 1) memory mosaics v2 consistently improve classification performance as it sees more demonstration shots (blue curves). In contrast, transformers struggle to maintain their performance and exhibit counterintuitively degraded performance as more demonstrations are provided (red curves). 2) Memory mosaics v2 significantly outperform transformers by more than 10%. Check Appendix F for the comparison between memory mosaics v2 small and transformer small (a larger margin).
@@ 6 Risk-return trade-off of frontier-model-sized memory mosaics v2
In contrast, taking the latter approach may require optimizing software and hardware, adapting techniques, a sharp sense of research direction, and possessing a keen sense of research direction along with strong problem solving abilities.131313 These requirements, in turn, demand a small group of high-quality researchers and managers. Despite the high requirements for personnel, this approach holds the potential for tremendous breakthroughs.
@@ 6.1 Comparison of two approaches
Table 5 shows the comparison on the new-knowledge storage and retrieval ability. Training on the same number of tokens (1t), transformers lag behind memory mosaics v2 by 12.3% (41.1% vs 53.4%). ×$\times 8$ times more training tokens (8t) improves the performance of transformers. However, the resulting transformer (trained on 8t tokens) still lags behind memory mosaics v2 (trained on 1t tokens) by 6.5% (46.9% vs 53.4%).
@@
Figure 5: Semantic label in-context learning comparison between memory mosaics v2 and transformer. Memory mosaics v2 is trained on 1t tokens, while three transformers are trained on 200b, 1t, 8t tokens, respectively. Transformer with ×$\times 8$ times more training data (8t, dash red line) starts to match the performance of Memory Mosaics v2 (1t, solid blue line).
@@
Figures 5 and 6 show the comparison on in-context learning ability. For semantic label tasks (Figure 5), ×$\times 8$ times more training data helps transformers (8t data) match the performance of memory mosaics v2 (1t data). However, for the more challenging anonymous label tasks, more training data cannot help transformers. Contour-intuitively, transformers trained on more training data (8t) exhibit a degraded performance on anonymous label tasks (Figure 6).
In summary, ×$\times 8$ more training data helps transformers in certain new task learning benchmarks. However, the resulting transformers (8t data) still lag behind memory mosaics v2 trained on 1t data. More importantly, in anonymous label tasks that heavily rely on the new task learning ability, more training data cannot help transformers. These experiments answer the initial question: “How much data does the transformer recipe approach need to match the performance of memory mosaics v2?”.
@@ Acknowledgments
Léon Bottou is a CIFAR fellow. We thank Gabriel Synnaeve, Jade Copet, Badr Youbi Idrissi, and Ammar Rizvi for their support with hardware, software, data, and baselines.
@@
Appendix A Gated time-variant key feature extractor & convocational value extractor
@@
Figure 7: Left: Key $k_{t}$ feature extractor: $k_{t}=\text{Norm}(\bar{k}_{t})$. Right: Value $v_{t}$ feature: $v_{t}=\text{Norm}(\bar{v}_{t})$.
@@ Appendix B Training data sequence length distributions
Appendix C Additional results for persistent-knowledge storage and retrieval
@@
Figure 6 shows six language benchmarks in which removing long-term memory from memory mosaics v2 after training degrades its performance.
@@
Appendix D Training details
hyperparameters
For all Memory Mosaics v2 and baseline Transformer models151515 The hidden dimension of the two-layers neural network in persistent memory is set to 6,144 and 14,336 for small and large models, respectively., we use a consistent set of hyperparameters. That is, a batch size of 1024, a sequence length of 4096, an adamw optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ accompanied by a $L_{2}$ weight decay of 0.1 and a gradient norm clip of 1, a learning rate warm-up of 2000 iterations followed by a cosine learning rate scheduler that reduces the learning rate by a factor of 100 at the end. The initial learning rates (after warm-up) are set to 3e-4 for “small” models and 1e-3 for “large” models.
Figure 9: Randomly overlapped long-term & short-term memory during training.
Appendix E Additional results for new-knowledge storage and retrieval
Table 7 shows that removing long-term memory from memory mosaics v2 after training degrades the performance on the ruler question-answer tasks by 20$\sim$30%. This indicates that the ruler question-answer tasks rely on long-term memory to perform well.
@@ Appendix E Additional results for new-knowledge storage and retrieval
Table 8: Comparison of Memory Mosaics v2 large (base model) and other public base models (similar scale) on ruler question-answer tasks. Memory Mosaics v2 large outperforms these models across all task lengths. The numbers in “*” rows come from Hsieh et al. [2024].
@@
Appendix F Additional results for in-context learning
Figure 10 and 11 shows the in-context learning comparison between memory mosaics v2 small and transformer small (llama-1.5b scale).
@@ Appendix F Additional results for in-context learning
Appendix G Prompt examples of multiclass classification tasks
G.1 Banking77 classification with semantic labels
@@
G.2 Goemotion classification with semantic labels
@@
G.3 Tacred classification with semantic labels
@@
Appendix H Computation resources

### v3

@@ 5.3 In-context learning
[…]
@@
Figure 3 compares the performance of memory mosaics v2 and transformers in three classification tasks with semantic target labels. The horizontal axis represents the number of shots, while the vertical axis represents the classification accuracy on $x_{test}$. We can observe two phenomena: 1) memory mosaics v2 consistently improve classification performance as it sees more demonstration shots (blue curves). In contrast, transformers struggle to maintain their performance and exhibit counterintuitively degraded performance as more demonstrations are provided (red curves). 2) Memory mosaics v2 significantly outperform transformers by more than 10%. Appendix H provides a similar comparison on a smaller model size ($\sim$1.5b), with an even larger margin. Appendix E further summarizes the comparison under matched model size or computation (FLOPs).
@@
Figure 5: Augmenting transformer with long-short term attention doesn’t help in-context learning.
Augment transformer with long-short term attention
Memory mosaics (v2) contains several unique components that are not applicable to transformers, such as the symmetric key and query, and the adaptive bandwidth. One seemingly applicable component for transformers is the separation of long-term and short-term memories introduced in Section 3.3. However, Figure 5 shows that augmenting a transformer with long-short-term attention does not help it overcome the limitations of in-context learning. These phenomena imply that memory mosaics (v2) is not simply a transformer variation but represents a different architecture.
Computation and Parameter Concerns
On the last two evaluation dimensions (new knowledge storage and retrieval, and in-context learning), memory mosaics v2 outperform transformers by more than 10% with slightly more parameters. This 10% advantage holds even when comparing under the same number of parameters or the same computational budget. See Appendix Figure 12 for details.
@@ 6 Risk-return trade-off of frontier-model-sized memory mosaics v2
In contrast, taking the latter approach may require optimizing software and hardware, adapting techniques, a sharp sense of research direction, and possessing a keen sense of research direction along with strong problem solving abilities. [^14] These requirements, in turn, demand a small group of high-quality researchers and managers. Despite the high requirements for personnel, this approach holds the potential for tremendous breakthroughs.
@@ 6.1 Comparison of two approaches
Table 5 shows the comparison on the new-knowledge storage and retrieval ability. Training on the same number of tokens (1t), transformers lag behind memory mosaics v2 by 12.3% (41.1% vs 53.4%). $\times 8$ times more training tokens (8t) improves the performance of transformers. However, the resulting transformer (trained on 8t tokens) still lags behind memory mosaics v2 (trained on 1t tokens) by 6.5% (46.9% vs 53.4%).
@@
Figure 6: Semantic label in-context learning comparison between memory mosaics v2 and transformer. Memory mosaics v2 is trained on 1t tokens, while three transformers are trained on 200b, 1t, 8t tokens, respectively. Transformer with $\times 8$ times more training data (8t, dash red line) starts to match the performance of Memory Mosaics v2 (1t, solid blue line).
@@
Figures 6 and 7 show the comparison on in-context learning ability. For semantic label tasks (Figure 6), $\times 8$ times more training data helps transformers (8t data) match the performance of memory mosaics v2 (1t data). However, for the more challenging anonymous label tasks, more training data cannot help transformers. Contour-intuitively, transformers trained on more training data (8t) exhibit a degraded performance on anonymous label tasks (Figure 7).
In summary, $\times 8$ more training data helps transformers in certain new task learning benchmarks. However, the resulting transformers (8t data) still lag behind memory mosaics v2 trained on 1t data. More importantly, in anonymous label tasks that heavily rely on the new task learning ability, more training data cannot help transformers. These experiments answer the initial question: “How much data does the transformer recipe approach need to match the performance of memory mosaics v2?”.
7 Fine-tuning speed: who can fine-tune with one minibatch?
Despite the strong in-context learning capability of memory mosaics v2 shown in Section 5.3, it may still be attractive to fine-tune a model for a specific domain in order to either reduce inference costs or improve in-domain performance. It is generally expected that such models can be efficiently fine-tuned for a new domain using a comparatively small number of examples.
Figure 8: Fine-tuning speed comparison between memory mosaics v2 and transformer.
Figure 8 compares the fine-tuning speed (in terms of data size) of memory mosaics v2 and transformers. Both models, pre-trained on 4k context windows, were fine-tuned to 32k context length using the recipe described in Section 4 and evaluated on the same ruler tasks (32k task-length) described in Section 5.2.
Surprisingly, a single fine-tuning mini-batch (one optimization step) on memory mosaics v2 yields a 22% accuracy improvement. Two fine-tuning mini-batches on memory mosaics v2 are sufficient to reach the optimal performance. In contrast, a transformer fine-tuned with 800 mini-batches still lags behind memory mosaics v2 fine-tuned with a single mini-batch.
@@ Acknowledgments
Léon Bottou is a CIFAR fellow. We thank Gabriel Synnaeve, Jade Copet, Badr Youbi Idrissi, and Ammar Rizvi for their considerable support with hardware, software, data, and baselines.
@@
Appendix A Gated time-variant key feature extractor & convolutional value extractor
@@
Figure 9: Left: Key $k_{t}$ feature extractor: $k_{t}=\text{Norm}(\bar{k}_{t})$. Right: Value $v_{t}$ feature extractor: $v_{t}=\alpha_{\psi}\text{Norm}(\bar{v}_{t})$.
@@ Appendix B Training data sequence length distributions
Appendix C Training details
hyperparameters
For all Memory Mosaics v2 and baseline Transformer models [^16] The hidden dimension of the two-layers neural network in persistent memory is set to 6,144 and 14,336 for small and large models, respectively., we use a consistent set of hyperparameters. That is, a batch size of 1024, a sequence length of 4096, an adamw optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.95$ accompanied by a $L_{2}$ weight decay of 0.1 and a gradient norm clip of 1, a learning rate warm-up of 2000 iterations followed by a cosine learning rate scheduler that reduces the learning rate by a factor of 100 at the end. The initial learning rates (after warm-up) are set to 3e-4 for “small” models and 1e-3 for “large” models.
Figure 11: Randomly overlapped long-term & short-term memory.
Parameter Initialization and reparameterization
Table 6 summarizes the parameter initialization methods and reparameterization tricks. $W_{1},W_{2},W_{3}$ refer to the parameters in persistent memory that are implemented as two-layer dense neural networks, $W_{2}\big(SiLU(W_{1}(x))*W3(x)\big)$. $SiLU(x)=x\cdot sigmoid(x)$ is an activation function. $d\in\{2048,4096\}$ indicates the hidden dimension of Memory Mosaics v2 small and large. $d^{\prime}\in\{6144,14336\}$ indicates the hidden dimension of the two-layer neural networks in persistent memory. $l$ indicates the depth of the Mosaics blocks, starting from 0.
Table 6: Parameter initialization methods and reparameterization tricks used in Memory Mosaics v2.
Parameter Location reparameterization Initialization
$\beta_{0}$ adaptive bandwidth $\beta_{0}=e^{\min(\theta,10)}$ $\theta=1.5$
$\beta_{1}$ adaptive bandwidth $\beta_{1}=e^{\min(\theta,10)}$ $\theta=1.5$
$\alpha$ adaptive bandwidth $\alpha=\min(|\theta|,1)$ $\theta=1/3$
$\alpha_{\psi}$ feature extractor $\alpha_{\psi}=e^{\min(|\theta|,15)}$ $\theta=0$
$\gamma$ feature extractor - $U(0,1)$
$W_{\psi},W_{\varphi},W_{g},W_{\lambda},W_{o}$ long-short memory - $\min\big(\max(\mathcal{N}(0,\sigma),-3\sigma),3\sigma\big)$, $\sigma=\frac{1}{\sqrt{2d(l+1)}}$
$W_{1},W_{3}$ persistent memory - $\min\big(\max(\mathcal{N}(0,\sigma),-3\sigma),3\sigma\big)$, $\sigma=\frac{1}{\sqrt{2d(l+1)}}$
$W_{2}$ persistent memory - $\min\big(\max(\mathcal{N}(0,\sigma),-3\sigma),3\sigma\big)$, $\sigma=\frac{1}{\sqrt{2d^{\prime}(l+1)}}$
$W_{e},W_{c}$ embedding & classifier - $\min\big(\max(\mathcal{N}(0,\sigma),-3\sigma),3\sigma\big)$, $\sigma=\frac{1}{\sqrt{2d}}$
Appendix D Failures of memory compression baselines
Many memory compression algorithms, such as rnns, xlstm [Beck et al., 2025], rwkv [Peng et al., 2023], and state-space models [Gu and Dao, 2023], fail on new-task storage and retrieval and in-context learning evaluation dimensions by construction. The reason is that these memory compression algorithms lack the ability to store large amounts of information before getting a command on how to process the information. One might argue to play around this shortage by reading the “command” before storing the large amounts of information. However, this process involves task-specific priori knowledge from human designers. In the end, instead of proving the machine is intelligent, it often proves that human designers are intelligent. Please recall that a child does not prepare all questions before going to school.
This incompetent of memory compression algorithms has been experimentally demonstrated by Hsieh et al. [2024] and Li et al. [2024] on both ruler benchmarks and in-context learning tasks.
Table 7 compares memory compression methods (rwkv-v5-7b and mamba-2.8b-slimpj) and non-compression method (llama2-7b) on ruler long-context tasks. It is clear that memory compression methods perform poorly as the required context length (i.e., required information storage space) increases.
Similarly, Table 8 compares memory compression methods (rwkv-5-world 7b and Mamba-2.8B) and non-compression method (qwen-1.5-7b-base and mistral-7b-v0.2-base) on in-context learning tasks (Tacred few-shot classification [Zhang et al., 2017]). In this challenging in-context scenario, memory compression methods just don’t work at all.
Please note that this section shouldn’t be used to criticize or hinder the study of memory compression methods. Memory compression methods have their advantages. In persistent-knowledge storage and retrieval evaluation dimension, they performs very well. For model efficiency, memory compression methods reveal a charming computation complexity. The goal of this section is to explain why this paper doesn’t choose memory compression methods as baselines.
Table 7: Comparison of memory compression methods (rwkv-v5-7b and mamba-2.8b-slimpj) and non-compression method (llama2-7b) on ruler long-context tasks. memory compression methods perform poorly as the required context length increases. Numbers are copied from Hsieh et al. [2024] Figure 4.
model task-length 1k task-length 2k task-length 4k
llama2-7b 96.0 91.6 95.0
rwkv-v5-7b 87.5 73.7 51.4
mamba-2.8b-slimpj 62.6 52.6 -
Table 8: Comparison of memory compression methods (rwkv-5-world 7b and Mamba-2.8B) and non-compression method (qwen-1.5-7b-base and mistral-7b-v0.2-base) on in-context learning tasks (Tacred few-shot classification [Zhang et al., 2017]). Memory compression methods fail on all cases. Numbers are copied from Li et al. [2024] Table 4.
model 1-shot 2-shots 3-shots 4-shots 5-shots
qwen-1.5-7b-base 7b 38.7 47.3 45.2 43.6 40.6
mistral-7b-v0.2-base 53.3 53.1 51.6 48.0 42.3
rwkv-5-world 7b 2.3 2.6 1.0 0 1.2
Mamba-2.8B 0 0 0 0 0
Appendix E Model efficiency comparison
As we emphasized in main text, model efficiency (e.g. model service, throughput, vram, etc.) is not the goal of this work. Many engineering works can be performed to adapt memory mosaics v2 to a custom use case or hardware. To aid in these potential adaptations, Figure 12 provides a model efficiency comparison in both computation (FLOPs) and model size (number of parameters) viewpoints. The results show that memory mosaics v2 outperforms transformer by more than 10% under either the same FLOPs or the parameters budgets.
Figure 12: Model efficiency (FLOPs or # parameters) comparison between Transformer large and Memory Mosaics v2 large. Top row shows the comparison on ruler “multi-unrelated-documents storing and question-answering” tasks, while the bottom row shows comparison on the Tacred in-context learning task. Memory mosaics v2 large outperforms transformer large by more than 10%.
Appendix F Additional results on persistent-knowledge storage and retrieval
@@
Table 9 shows six language benchmarks in which removing long-term memory from memory mosaics v2 after training degrades its performance.
@@
Appendix G Additional results on new-knowledge storage and retrieval
[…]
Table 12 illustrates the effect of the stochastic long-term memory size training setup introduced in Section C. This stochastic long-term memory size setup is used to encourage the allocation of position-invariant signals and position-dependent signals to long-term and short-term memories.
Table 13 compares memory mosaics v2 and transformers on a typical ‘needle-in-a-haystack’ task from ruler [Hsieh et al., 2024]. The typical ‘needle-in-a-haystack’ is too easy such that many models can achieve a near-perfect performance.
@@ Appendix G Additional results on new-knowledge storage and retrieval
Table 11: Comparison of Memory Mosaics v2 large (base model) and other public base models (similar scale) on ruler question-answer tasks. Memory Mosaics v2 large outperforms these models across all task lengths, despite that Memory Mosaics v2 uses 1/4 generation lengths (32 tokens) of other public base models (128 tokens). The numbers in “*” rows come from Hsieh et al. [2024].
@@
Table 12: The effect of stochastic long-term memory size (during training) in memory mosaics v2 small model on ruler question-answer tasks. Both models are trained on 4k context length, then evaluated on 32k context length without any fine-tuning. The stochastic long-term memory size setup boost context length extrapolation ability by more than 15%.
model context length stochastic long-term memory task-length 4k task-length 32k
memory mosaics v2 small 4k No 43.6 15.9
memory mosaics v2 small 4k Yes 45.0 31.7 (+15.8)
Table 13: ruler s-niah benchmark comparison between transformer and memory mosaics v2.
model context length 4k 8k 16k task-length 32k
transformer small 32k 99.4 99.0 98.2 97.8
memory mosaics v2 small 32k 100.0 100.0 100.0 100.0
transformer large 32k 100.0 100.0 100.0 99.6
memory mosaics v2 large 32k 100.0 100.0 100.0 100.0
Appendix H Additional results for in-context learning
Figure 13 and 14 shows the in-context learning comparison between memory mosaics v2 small and transformer small (llama-1.5b scale).
@@ Appendix H Additional results for in-context learning
Appendix I Prompt examples of multiclass classification tasks
I.1 Banking77 classification with semantic labels
@@
For each prompt with either “[return]” or “[space]” delimiter, we also try to shuffle the demonstration example (i.e., service query: […], intent category:[…]) orders within each one shot. This shuffling process provides another two more prompts.
I.2 Banking77 classification with anonymous labels
Anonymous tasks use the same set of prompts except that anonymous tasks replace semantic labels (e.g. city_arrival, city_linking) with anonymous labels (e.g. class_00, class_01).
I.3 Goemotion classification with semantic labels
@@
For each prompt with either “[return]” or “[space]” delimiter, we also try to shuffle the demonstration example orders within each one shot. This shuffling process provides another two more prompts.
I.4 Goemotion classification with anonymous labels
Anonymous tasks use the same set of prompts except that anonymous tasks replace semantic labels with anonymous labels (e.g. class_00, class_01).
I.5 Tacred classification with semantic labels
@@
For each prompt with either “[return]” or “[space]” delimiter, we also try to shuffle the demonstration example orders within each one shot. This shuffling process provides another two more prompts.
I.6 Tacred classification with anonymous labels
Anonymous tasks use the same set of prompts except that anonymous tasks replace semantic labels with anonymous labels (e.g. class_00, class_01).
Appendix J Computation resources